### v1

# Adaptive Vision-Guided Robotic Arm Control for Precision Pruning in Dynamic Orchard Environments

Abstract
This study presents a vision-guided robotic control system for automated fruit tree pruning applications. Traditional agricultural practices rely on labor-intensive tasks and processes that lack scalability and efficiency, creating a pressing need for automation research to address growing demands for higher crop yields, scalable operations, and reduced manual labor. To this end, this paper proposes a novel algorithm for robust and automated fruit pruning in dense orchards. The proposed algorithm utilizes CoTracker, that is designed to track 2D feature points in video sequences with significant robustness and accuracy, while leveraging joint attention mechanisms to account for inter-point dependencies, enabling robust and precise tracking under challenging and sophisticated conditions. To validate the efficacy of CoTracker, a Universal Robots manipulator UR5e is employed in a Gazebo simulation environment mounted on ClearPath Robotics Warthog robot featuring an Intel RealSense D435 camera. The system achieved a 93% success rate in pruning trials and with an average end trajectory error of 0.23 mm. The vision controller demonstrated robust performance in handling occlusions and maintaining stable trajectories as the arm move towards the target point. The results validate the effectiveness of integrating vision-based tracking with kinematic control for precision agricultural tasks. Future work will focus on real-world implementation and the integration of 3D reconstruction techniques for enhanced adaptability in dynamic environments.
keywords:
visual servoing, agricultural automation, robotic manipulator, perception and sensing, agricultural robotics, precision pruning.
thanks: submitted to IFAC AgriControl 2025
1 Introduction
The rapidly growing population and urbanization are driving a sharp increase in food demand. Agriculture remains essential for producing fruits and crops to meet this need. However, modern agricultural practices still rely heavily on human labor, limiting crop yield, scalability, and efficiency. Pruning, a vital agricultural task, improves plant health, fruit production, and overall quality but remains labor-intensive, requiring expertise and significant manual effort (Zhao et al., 2016). With a severe shortage of skilled labor and high costs associated with manual pruning, automation has become a key research focus (Fimiani et al., 2023). Recent efforts have explored robotic pruning systems with advanced vision and control algorithms, offering a sustainable, precision-based alternative to manual labor.
Despite significant progress in many areas of agricultural robotics (e.g., harvesting, thinning, and crop scouting), pruning still remains a formidable challenge. First, fruit trees exhibit highly unstructured, dynamic environments with complex branch geometries and frequent occlusions by other branches or orchard infrastructure. Second, precise manipulation is required to position a cutting tool at the exact pruning point and orientation while avoiding collisions and excessive forces. Therefore, there is a pressing need for accurate perception, robust point tracking, and agile control schemes to enhance the capability and practical adaptability of robotic pruning systems.
Figure 1: Simulation environment showing the Gazebo setup with the dormant apple trees and a Warthog bot with a mounted UR5e arm.
Automated pruning systems require advanced sensing, modeling, and control to function effectively in dynamic orchard environments. The process begins with accurately sensing and modeling tree structures using sensors like stereo vision or LiDAR, combined with 3D point-cloud registration, to capture branch geometry and spatial arrangement (Botterill et al., 2017; Tabb and Medeiros, 2017). Our previous research demonstrated that accurate branch diameter estimation (Ahmed et al., 2025), along with bud detection and counting (Ahmed et al., 2024), provides essential data for determining optimal crop loads. These measurements inform data-driven pruning rules to identify branches and pruning locations while balancing tree structure maintenance and fruit production. After selecting pruning points, a precise, collision-free trajectory must be planned to guide the cutting tool while avoiding obstacles such as trunks, wires, and posts. However, traditional open-loop or purely position-based methods can accumulate significant errors in real-world conditions, reducing accuracy and precision (You et al., 2020, 2023).
Recent advancements in visual servoing address operational challenges in robotic pruning through closed-loop control based on camera feedback (Shamshiri et al., 2023; Dong and Zhu, 2015). While position-based visual servoing (PBVS) relies on 3D pose estimation, image-based approaches (IBVS) minimize 2D feature errors, offering greater robustness in unstructured orchards where branches, foliage, and infrastructure complicate calibration. Due to its accuracy and precision, visual servoing has been widely explored for tool alignment in dynamic pruning environments. Yandun et al. (Yandun et al., 2021) used deep reinforcement learning (DRL) trained on 3D vine models to navigate cluttered canopies, demonstrating adaptability but requiring extensive training. You et al. (2022) improved this by introducing a hybrid vision/force control framework, where vision-based policies trained on synthetic data guided a cutter to sub-centimeter accuracy, with force feedback mitigating excessive contact with rigid branches. Gebrayel et al. (2024) refined PBVS using iterative closest point (ICP) variants for real-time vine alignment but struggled in highly occluded scenes due to reliance on continuous point-cloud tracking. While RL enables collision avoidance, its dependence on simulated training limits robustness to occlusions, and wind-induced vine motion remains a challenge. These limitations highlight the need for robust visual tracking methods that maintain accuracy despite dense foliage and occlusions.
To overcome orchard challenges such as occlusions, wind-induced movements (Spatz and Theckes, 2013), and variable lighting, tracking methods have advanced beyond keypoint detection and optical flow. These newer approaches handle repetitive textures, occlusions, and rapid motion shifts common in agriculture (Matos et al., 2024). Recent developments include LoCoTrack (Cho et al., 2024), which enhances robustness with 4D correlation volumes for arbitrary point tracking. Tapir (Doersch et al., 2023) employs transformer-based global matching for long-term tracking but remains computationally demanding and struggles in ambiguous regions. CoTracker3 (Karaev et al., 2024a) addresses these limitations with a streamlined architecture that replaces heavy correlation processing with lightweight MLPs, achieving 27% faster performance than LoCoTrack while maintaining accuracy. Its pseudo-labeling approach allows training on real-world videos without manual annotations, improving generalization and reducing data requirements. These features make CoTracker3 well-suited for dynamic orchard environments requiring both precision and efficiency.
This research aims to develop an autonomous robotic system for apple orchard pruning. The robot navigates orchard rows independently, analyzing dormant trees to assess branch geometry, orientation, and bud distribution. Using this structural data, it identifies optimal pruning locations, tracks cutting points with a transformer-based model, and executes precise cuts with a UR5e robotic arm and a cutting tool. To advance autonomous pruning, this work presents an adaptive system that integrates CoTracker3’s robust point-tracking with differential inverse kinematics (IK) and proportional velocity control. Operating primarily in the 2D image plane, the system continuously updates motion based on feature positions, ensuring precise and responsive control. The approach is validated in a Gazebo simulation using a UR5e robotic arm (Universal Robots, Denmark) and an Intel RealSense D435 camera (Intel, California, US).
This paper is organized as follows. Section 2 outlines methodological details including a brief description of CoTracker3. Section 3 presents key results and provides a comprehensive discussion. Finally, Section 4 provides a brief conclusion along with some remarks on future work.
2 Methodology
Figure 2: System Architecture (Karaev et al., 2024a): The tracking system uses a camera mounted on the robotic arm to capture visual data and feed to Cotracker3’s perception pipeline to extract convolutional features from each frame. The system analyzes feature correlations between frames to track the pruning point. A transformer iteratively refines the pruning point’s track, confidence, and visibility using previous estimates for accurate tracking.
This section outlines our approach to precision pruning of apple trees. We use the Gazebo simulation platform, integrated with the Robot Operating System (ROS), to create a realistic testing environment, as shown in Figure 1. The robotic system consists of a Universal Robots UR5e manipulator with six degrees of freedom and an Intel RealSense D435 camera mounted on its end-effector. The system’s primary goal is to visually servo the robotic arm toward the pruning branch, ensuring precise end-effector alignment with the target pruning point.
Gazebo serves as the core simulation environment, providing a physics-based framework to model apple trees, the robotic manipulator, and the surrounding workspace. It enables accurate rendering of object interactions, collision dynamics, and sensor feedback, closely replicating real-world pruning conditions. Widely used in robotics research, Gazebo offers a high-fidelity physics engine, sensor integration, and seamless compatibility with ROS, facilitating efficient testing before real-world deployment (Koenig and Howard, 2004).
The vision controller utilizes CoTracker3 to track pruning points across video frames. Currently, pruning points are selected manually, but ongoing research explores AI-driven methods for automatic pruning point detection based on tree geometry. These advancements will be integrated into the system in future iterations. By operating directly in image space rather than reconstructing 3D depth data, the system maintains a streamlined visual servoing approach. The following sections detail the key components of our methodology, including control architecture, feature tracking, controller implementation, and performance evaluation.
2.1 CoTracker3
CoTracker3 is a transformer-based point tracking model that builds on previous work in CoTracker while simplifying the architecture and improving tracking performance (Karaev et al., 2024b). It is designed for robust tracking of 2D points in video sequences. Unlike traditional methods that track points independently, CoTracker3 leverages joint attention mechanisms to track multiple points jointly, enabling superior tracking performance even under challenging conditions such as occlusions or when points move out of the field of view. This joint tracking approach allows the model to infer the positions of occluded points using information from visible points, significantly improving tracking robustness. By operating on overlapping sliding windows of image frames, CoTracker3 ensures continuity and accuracy over long video sequences.
The CoTracker3 model begins with the user selecting an initial point of interest $(x_{0},y_{0})$ in the camera frame, which serves as the query embedding $\mathbf{q}$. This query embedding initializes the tracking process. CoTracker3 updates the position of the point dynamically across frames by attending to temporal and spatial features in the sliding window of video frames. The tracking operation is formally described as:
$(x_{t},y_{t})=\mathcal{T}\left((x_{t-1},y_{t-1}),\mathbf{F}_{t-1:t},\mathbf{q}\right),$ (1)
where $\mathcal{T}$ represents the CoTracker3 function, $\mathbf{F}_{t-1:t}$ encapsulates the feature embeddings from the sliding window of frames $t-1$ to $t$, and $\mathbf{q}$ is the query embedding derived from the user-selected point $(x_{0},y_{0})$. The model refines the position $(x_{t},y_{t})$ iteratively by leveraging attention mechanisms that link the spatial and temporal contexts. Branch occlusion is a significant challenge in tree pruning scenarios due to the cluttered environment. When a tracked point becomes occluded by branches, CoTracker’s joint attention capabilities infer the most probable position based on the contextual information from visible points and prior frames. This mechanism allows the model to maintain continuity in tracking until the point reappears. To evaluate the performance of the tracking, CoTracker3 computes an error vector between the tracked point $(x_{t},y_{t})$ and the center of the image frame $(x_{c},y_{c})$:
$\mathbf{e}_{t}=\begin{bmatrix}e_{x}\\ e_{y}\end{bmatrix}=\begin{bmatrix}x_{c}-x_{t}\\ y_{c}-y_{t}\end{bmatrix}.$ (2)
The error vector $\mathbf{e}_{t}$ is used to guide the robotic controller in minimizing positional deviations by adjusting the manipulator’s position accordingly. The error vector’s magnitude, $\|\mathbf{e}_{t}\|$, serves as a key metric for assessing tracking accuracy and guiding corrective actions. By combining these mechanisms, CoTracker3 provides a robust and efficient solution for maintaining the pruning point’s position in the camera frame, ensuring seamless and accurate robotic manipulation even in complex environments with branch occlusions.
2.2 Low Level Joint Controller
We adopt a proportional (P) control strategy that operates directly on the 2D image-space error provided by the vision module (CoTracker3). Let $(x_{c},y_{c})$ denote the image center and $(x_{p},y_{p})$ the coordinates of the pruning point as tracked in the camera feed. The error vector is defined as
$\mathbf{e}\;=\;\begin{bmatrix}e_{x}\\ e_{y}\end{bmatrix}\;=\;\begin{bmatrix}x_{c}-x_{p}\\ y_{c}-y_{p}\end{bmatrix}.$ (3)
Denoting proportional gains by $K_{p}^{x}$ and $K_{p}^{y}$, the incremental motion commands in the image plane are calculated as
$\Delta x\;=\;K_{p}^{x}\cdot e_{x},\quad\Delta y\;=\;K_{p}^{y}\cdot e_{y}.$ (4)
These planar commands define how the end-effector should move in the camera frame to reduce the 2D error. However, the manipulator itself must execute these displacements through joint-space motion. The next subsection details how we map these commands to corresponding joint angle updates via a differential inverse kinematics (IK) approach.
2.3 Differential Inverse Kinematics
To translate the image-based control signals into robot joint motions, we employ a differential inverse kinematics (IK) approach augmented by a deliberate forward step size along the end-effector $z$-axis. This design allows the robot to continuously advance towards the branch while correcting its $x$$y$ alignment based on CoTracker3 feedback.
First, we define the manipulator’s kinematics using the Denavit–Hartenberg (DH) convention. The forward kinematics map joint angles $\mathbf{q}\in\mathbb{R}^{6}$ to a transformation matrix $\mathbf{T}_{\mathrm{end}}(\mathbf{q})\in\mathbb{R}^{4\times 4}$, which encodes the end-effector pose (position and orientation) in 3D space. Denoting by $\mathbf{J}(\mathbf{q})\in\mathbb{R}^{6\times 6}$ the Jacobian of $\mathbf{T}_{\mathrm{end}}$ w.r.t. the joint angles, we have
$\dot{\mathbf{x}}_{\mathrm{end}}\;=\;\mathbf{J}(\mathbf{q})\,\dot{\mathbf{q}},$ (5)
where $\dot{\mathbf{x}}_{\mathrm{end}}\in\mathbb{R}^{6}$ is the spatial velocity of the end-effector (three components for translation, three for rotation), and $\dot{\mathbf{q}}\in\mathbb{R}^{6}$ is the vector of joint velocities.
At each control loop, the vision module (CoTracker3 plus a proportional controller) supplies horizontal and vertical offsets $(\Delta x,\Delta y)$ in image space. We map these offsets to a small planar motion command in the manipulator frame while also prescribing a constant incremental step $\Delta z>0$ along the end-effector’s $z$-axis to move the robot steadily forward toward the pruning target. Symbolically, we assemble a 6D velocity vector:
$\dot{\mathbf{x}}_{\mathrm{end}}\;=\;\begin{bmatrix}\Delta x,\Delta y,\Delta z,0,0,0\end{bmatrix}^{\top},$ (6)
where $\Delta x$ and $\Delta y$ reflect the image-space alignment corrections (appropriately scaled into physical units), and $\Delta z$ is the chosen forward step size in meters per control step (remaining zeros in the equation are for the rotational components).
To realize these Cartesian velocities via joint-space commands, we invert the Jacobian:
$\dot{\mathbf{q}}\;=\;\mathbf{J}^{+}(\mathbf{q})\,\dot{\mathbf{x}}_{\mathrm{end}},$ (7)
where $\mathbf{J}^{+}$ denotes the pseudo-inverse of the Jacobian. Integrating $\dot{\mathbf{q}}$ over a short timestep yields the new joint angles $\mathbf{q}\leftarrow\mathbf{q}+\dot{\mathbf{q}}\,\Delta t$. This iterative scheme continues until the image-space error $(x_{c}-x_{p},\,y_{c}-y_{p})$ remains within a desired tolerance, indicating that the end-effector stays visually locked on the pruning point while advancing forward.
In practice, gains can be tuned to ensure smooth motion, and the forward step size $\Delta z$ can be adjusted according to task needs (e.g., slow approach for delicate cuts). By combining vision-based $x$$y$ corrections with a steady $z$-axis motion in the differential IK framework, the manipulator effectively zeros out lateral errors while proceeding directly toward the pruning location. This integrated control approach yields robust pruning trajectories even under moderate occlusions or shifting tree branches due to wind.
2.4 Evaluation Metrics
The performance of the vision-based controller was evaluated using several quantitative metrics. The trajectory error, $E_{traj}$, is calculated using desired and target joint angles which are then converted into cartesian space and error is computed using the following equation:
$E_{traj}=\sqrt{(x_{c}-x_{t})^{2}+(y_{c}-y_{t})^{2}+(z_{c}-z_{t})^{2}},$ (8)
where $(x_{c},y_{c},z_{c})$ denotes the current end-effector position, and $(x_{t},y_{t},z_{t})$ is the target trajectory location. The success rate, $S$, measures the proportion of trials in which the end-effector successfully aligned with the pruning point with a tolerance of 5 millimeter:
$S=\frac{\text{Number of Successful Trials}}{\text{Total Trials}}\times 100\%.$ (9)
3 Results and Discussion
Table 1: Summary of Quantitative Results. Key performance metrics showing accuracy, success rate, and response time from simulation trials with the vision-guided controller.
Metric Mean Value Description
End Trajectory Error (mm) 0.23 Error in positioning pruning point
Success Rate (%) 93 Percentage of successful trials
Response Time (s) 2.34 Time to align pruning point
This section presents the experimental results and analysis of the vision-based controller’s performance in apple tree pruning. Table 1 summarizes the results from 20 simulated pruning trials. To test the system’s robustness, the robot was placed in front of the tree trunk, and pruning points were selected at the farthest edges of branches. This setup required the controller to make significant directional adjustments. The system successfully completed 93% of the trials, with an average end trajectory error of 0.23 mm in successful attempts. Figure 3 shows the error trends over time during a typical pruning trial. At the start, the error magnitude decreased rapidly as the proportional controller adjusted the end-effector position. Further refinements using the differential inverse kinematics (IK) approach continued to reduce the error until it reached a steady-state value.
Figure 3: A plot showing the error between the image center and target point converging to zero with successive iterations of the algorithm over time.
The results validate the effectiveness of the proposed vision-based controller in handling precision pruning tasks. The integration of CoTracker3, proportional control, and differential IK enables accurate, reliable, and efficient pruning, even in complex environments. However, the trials also revealed several failure modes that need to be addressed for further improvement. One common failure occurred due to manipulator singularities, where certain joint configurations caused control instability. Some failures also happened when target points were beyond the robot’s kinematic reach. Another limitation was the system’s inability to replan trajectories when occlusions blocked the target, leading to task interruptions. Additionally, in dense branching areas, the tracking algorithm sometimes lost the original target and mistakenly switched to a visually similar branch nearby. These challenges indicate the need for enhancements in trajectory planning and target tracking to improve the system’s overall robustness and reliability in real-world pruning scenarios.
In our real-world dataset, collected from a dormant apple orchard, CoTracker3 demonstrated robust temporal tracking of pruning targets. As illustrated in Figure 4, a close-up view of four consecutive frames highlights how CoTracker3 consistently tracks distinct branch points over time. The model leveraged spatiotemporal attention, ensuring smooth tracking continuity as the camera moves. By incorporating learned feature embeddings, CoTracker3 avoided drift and maintained stable predictions for each tracked keypoint.
(a) Frame at $t_{1}$
(b) Frame at $t_{2}$
(c) Frame at $t_{3}$
(d) Frame at $t_{4}$
Figure 4: Tracking of points over four consecutive frames. Each subfigure represents a frame at time $t_{1},t_{2},t_{3},t_{4}$, showing the consistency of CoTracker3.
Beyond its ability to track points through time, CoTracker3 also provides valuable information into occlusion handling, a common challenge in orchard environments where overlapping branches often obscure pruning targets. As shown in Figure 5, CoTracker3 accurately tracked partially ocluded point. When a point became occluded, the model did not simply discard the tracking instance; rather, it predicted its likely location using information from adjacent frames and spatially correlated features. This ability was helpful to smoothly resume tracking once the point re-emerged into the scene. By using occlusion and continuity information, CoTracker3 provides a reliable mechanism for robotic pruning systems to make informed decisions about when a target is temporarily occluded. This capability is expected to substantially improve the robustness of autonomous pruning operations in dense orchard environments.
(a) Frame at $t_{1}$
(b) Frame at $t_{2}$
Figure 5: Occlusion handling in CoTracker3. a) frame $t_{1}$ , the target is fully visible. b) In frame $t_{2}$, the target becomes occluded by another branch, yet tracking remains consistent, demonstrating CoTracker3’s robustness in challenging orchard environments.
Qualitative observation across a set of dormant pruning images showed that the model demonstrated accurate tracking in most cases. However, a few challenging instances were observed. Specifically, the issue was observed when a target completely disappears from few frames due to changes in camera perspective. CoTracker3 occasionally struggled to re-identify previously tracked points when they reappeared. This behavior is attributed to the complexity of orchard environments, where branches often look similar, making re-initialization of lost points difficult. Despite these challenges, CoTracker3 demonstrated a 93% success rate in simulation, confirming its reliability for pruning tasks. For practical deployment, the system required a minimum frame rate of 15-20 Hz to effectively track branch movements. However, the current implementation in simulation operated at lower rates. Future work will focus on optimizing computational efficiency to meet real-time requirements for orchard applications.
4 Conclusion and Future Work
The contemporary agricultural practices rely heavily on manual and intensive human-labor centric processes reducing crop-yield and limiting scalability. Pruning is one of the agricultural techniques that to-date remains a manual process. To this end, this paper proposed an automation algorithm for manual labor-free fruit-tree pruning in dense orchards. The proposed framework utilized CoTracker3, a transformer based robust and real-time feature tracking algorithm, to track the points of interest on fruit-trees and guide a robotic manipulator based pruning tool to the pruning point under extremely challenging and occluded scenes. The framework consists of a feature tracking layer, employing CoTracker3, a planning layer, consisting of a differential inverse kinematics algorithm, and a low-level proportional controller. Extensive experimental and simulation validations demonstrate a success rate of 93% and average trajectory tracking error of 0.23 mm. Future work includes testing the system in actual orchards to assess its performance under varying environmental factors. Furthermore, integrating 3D reconstruction techniques and deformable branch modeling would allow for a more adaptive and precise control strategy, improving the system’s ability to handle dynamic and cluttered environments.
{ack}
This research is funded by the National Science Foundation and United States Department of Agriculture, National Institute of Food and Agriculture, United States through the following grants: AWD003473 as part of the “AI Institute for Agriculture (AgAID)” Program and 2023-67021-38908 under the International Collaboration Grant.